OpenStudy (anonymous):

Find the sum of the first six terms of the sequence for which a2 = 0.7 and a3 = 0.49

OpenStudy (anonymous):

@terenzreignz sorry! I just have one more

terenzreignz (terenzreignz):

These questions are so different from one another... this isn't an exam, is it? -_-

OpenStudy (anonymous):

No it is not an exam... it is a study guide for my test tomorrow

OpenStudy (campbell_st):

well the series is geometric based on the information a general term is \[a_{n} = a_{1} r^{n-1}\] a1 = 1st term, r = common ratio so start by finding the common ratio 2nd term \[0.7 = ar\] 3rd term \[0.49 = ar^2... or..... 0.49 = (ar) \times r ...or...... 0.49 = 0.7 \times r\] so the common ratio is 0.7 and the 1st term is 1 then the you can use this information with n = 6 to find the sum

OpenStudy (anonymous):

so it goes 1,0.7,0.49...

OpenStudy (anonymous):

So is 2.94 the sum of the first 6?